How to graph these systems of equations
a. 2x-5y = 11
3x+ y = 8
b. 4x+3y = 1
3x+2y = 2

Answer by Alan3354(69443)

How to graph these systems of equations
a. 2x-5y = 11
3x+ y = 8
b. 4x+3y = 1
3x+2y = 2
------------------
They're all straight lines, so you need 2 points.
---
2x-5y = 11
Pick 2 values for x and find y
x = 0; y = -11/5
x = 5; y = -1/5
Plot the points, draw a line thru them.
Do the others the same way.